# TabRep: Training Tabular Diffusion Models with a Simple and Effective Continuous Representation

### v1

@@ 1 Introduction
Since diffusion-based generative models rely on continuous transformations of denoising score-matching, or invertible mappings between data and latent spaces, designing an effective data representation is key for high-performing diffusion models [5]. In this work, we propose TabRep, a simple and effective continuous representation framework for training tabular diffusion models. We introduce geometric insights to understand the properties of a continuous data manifold for tabular diffusion models. Then, we craft an abstract and useful representation, that enables diffusion models like DDPM [20] and Flow Matching [31] to capture the posterior distribution efficaciously. TabRep’s representation maintains a dense two-dimensional representation, circumventing the curse of dimensionality. It is also well-suited to model the cardinality of nominal features, ensuring ample separability to distinguish between different categories. Lastly, coupling the cyclical nature of our representation with the separability it offers enables it to facilitate both ordinal and periodic features. These attributes present desirable characteristics that make it easier for diffusion models to extract meaningful information from a unified continuous tabular data representation. We conduct comprehensive experiments against state-of-the-art diffusion baselines across various datasets and benchmarks. The results showcase that TabRep consistently outperforms these baselines in quality, privacy preservation, and computation cost, indicating our superior capabilities to generate tabular data. Our architecture can be found in Figure 1.
@@ 4.1 Geometric Implications on the Data Representation
In traditional deep learning [16], a sparse representation, such as one-hot encoding, suffers from the curse of dimensionality [3], where the feature space grows exponentially with the number of categories, reducing model generalization [27, 41]. While maintaining a dense representation is key, representation learning also [4, 29] emphasizes the importance of separability, enabling the neural network to learn decision boundaries in the continuous embedding space. In the following excerpt, we find that balancing both density and separability is essential for unified tabular diffusion.
Density. Diffusion models, which learn to generate data through iterative perturbations and reconstructions of noise, encounter geometric challenges when applied to high-dimensional, sparse representations of categorical features. In this discussion, we use the one-hot representation to provide geometric insights into these challenges.
@@
Diffusion models rely on gradients derived from a learned vector field to denoise data iteratively [20, 31]. For regions within proximity of the singular hyperplanes, gradients suffer from high variance due to conflicting directions arising from equidistant one-hot points. We show that the variance of the gradients from a learned vector field increases asymptotically with respect to $n$-singular points. Assume $x$ is a noisy observation from a Gaussian centered at a weighted one-hot vector $\alpha_{t}e_{k}\in\mathbb{R}^{K}$. We can define the forward diffusion process as: $p_{t}(x|e_{k})=\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I)$. We derive the variance of learning gradients at an $n$-singular point as:
$\mathrm{Var}(g|x)=\frac{\alpha_{t}^{2}}{\sigma_{t}^{4}}\frac{n-1}{n\sigma_{t}^{4}},$ (10)
@@
where we define the gradient and expected gradient as $g$ and $\bar{g}$. See Appendix A.1 for a proof. We find that near an $n$-singular point, $x$, the posterior-weighted variance of the score function is strictly positive and increases asymptotically with $n$. In contrast, at a non-singular point, the posterior leans towards $e_{k^{*}}$, causing the variance to reduce to 00.
@@
Separability and Order. A sparse representation naturally accommodates nominal features, enabling one-hot to assign each category to each dimension. However, higher-dimensional spaces introduce higher-order singular hyperplanes that complicate training. This presents a density-separability trade-off. In our experiments (Table 3), we discover that sparse representations has a significant impact on harming diffusion generative performance. Thus, our initial goal is to reduce the dimensionality when designing our representation.
@@ 6 Conclusion
In this work, we present TabRep a tabular diffusion architecture trained with a simple yet effective unified continuous representation. Motivated by geometric implications on the data manifold, our representation is dense, separable, and captures meaningful information for diffusion models. We conducted extensive experiments on a myriad of datasets, baselines, and benchmarks to evaluate TabRep. The results showcase TabRep’s prowess in generating high-quality, privacy-preserving synthetic data while remaining computationally inexpensive.
@@ A.1 Variance of Learning Gradients in Diffusion Models
We show that the variance of the gradients from a learned vector field increases asymptotically with respect to $n$-singular points relative to non-singular points. Assume $x$ is a noisy observation from a Gaussian centered at a weighted one-hot vector $\alpha_{t}e_{k}\in\mathbb{R}^{K}$. We can define the forward diffusion process as:
@@
Hence, the score of the distribution is as follows:
@@
$q(e_{k}|x)$ $=\frac{q(x|e_{k})p(e_{k})}{\sum_{m=1}^{K}q(x|e_{m})p(e_{m})}$ (21)
$=\frac{q(x|e_{k})}{\sum_{m=1}^{K}q(x|e_{m})}$ (22)
$=\frac{q(x|e_{k})}{\sum_{k\in S}q(x|e_{k})+\sum_{m\notin S}q(x|e_{m})}$ (23)
Since the forward process is modeled as: $q(x|e_{k})=\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I)$, we can infer that:
$q(x|e_{k})\propto\exp\left(-\frac{1}{2\sigma_{t}^{2}}||x-\alpha_{t}e_{k}||^{2}\right)$ (24)

### v2

# TabRep: a Simple and Effective Continuous Representation for Training Tabular Diffusion Models

@@ 1 Introduction
Since diffusion-based generative models rely on continuous transformations of denoising score-matching, or invertible mappings between data and latent spaces, designing an effective data representation is key for high-performing diffusion models [5]. In this work, we propose TabRep, a simple and effective continuous representation for training tabular diffusion models. We introduce geometric insights to understand the properties of a continuous data manifold for tabular diffusion models. Then, we craft an abstract and useful representation, that enables diffusion models like DDPM [20] and Flow Matching [31] to capture the posterior distribution efficaciously. TabRep’s representation maintains a dense two-dimensional representation, circumventing the curse of dimensionality. It is also well-suited to model the cardinality of nominal features, ensuring ample separability to distinguish between different categories. Lastly, coupling the cyclical nature of our representation with the separability it offers enables it to facilitate ordinal features. These attributes present desirable characteristics that make it easier for diffusion models to extract meaningful information from a unified continuous tabular data representation. We conduct comprehensive experiments against tabular diffusion baselines across various datasets and benchmarks. The results showcase that TabRep consistently outperforms these baselines in quality, privacy preservation, and computation cost, indicating our superior capabilities to generate tabular data. Our architecture is in Figure 1.
@@ 4.1 Geometric Implications on the Data Representation
In traditional deep learning [16], a sparse representation suffers from the curse of dimensionality [3], where the feature space grows exponentially with the number of categories, reducing model generalization [27, 41]. While maintaining a dense representation is key, representation learning also [4, 29] emphasizes the importance of separability, enabling the neural network to learn decision boundaries in the continuous embedding space. In the following excerpt, we find that balancing both density and separability while incorporating order is essential for unified tabular diffusion.
Density. Diffusion models, which learn to generate data through iterative perturbations and reconstructions of noise, encounter geometric challenges when applied to high-dimensional, sparse representations of categorical features. In this discussion, we use the sparse one-hot representation as an exemplar to provide geometric insights into these challenges.
@@
Diffusion models rely on gradients derived from a learned vector field to denoise data iteratively [20, 31]. For regions within proximity of the singular hyperplanes, gradients suffer from high variance due to conflicting directions arising from equidistant one-hot points. We show that the variance of the score function increases asymptotically with the degree of $n$-singular points. Assume $x$ is a noisy observation from a Gaussian centered at a weighted one-hot vector $\alpha_{t}e_{k}\in\mathbb{R}^{K}$. We can define the forward diffusion process as: $p_{t}(x|e_{k})=\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I)$. We derive the variance of the conditional score function evaluated at an $n$-singular point as:
$\mathrm{Var}(g|x)=\frac{\alpha_{t}^{2}}{\sigma_{t}^{4}}\frac{n-1}{n},$ (10)
@@
where we define the conditional score and expected score as $g$ and $\bar{g}$. See Appendix A.1 for a proof. We find that near an $n$-singular point, $x$, the posterior-weighted variance of the conditional score function is strictly positive and increases asymptotically with $n$. In contrast, at a non-singular point, the posterior leans towards $e_{k^{*}}$, leading the score variance to approach zero.
@@
Separability and Order. A sparse representation naturally accommodates separability for nominal features, enabling one-hot to assign each category to each dimension. However, higher-dimensional spaces introduce higher-order singular hyperplanes that complicate training. This presents a density-separability trade-off. In our experiments (Table 3), we discover that sparse representations has a significant impact on harming diffusion generative performance. Thus, our initial goal is to reduce the dimensionality when designing our representation.
@@ 6 Conclusion
In this work, we present TabRep, a simple and effective continuous representation for training tabular diffusion models. Motivated by geometric implications on the data manifold, our representation is dense, separable, and captures meaningful information for diffusion models. We conducted extensive experiments on a myriad of datasets, baselines, and benchmarks to evaluate TabRep. The results showcase TabRep’s prowess in generating high-quality, privacy-preserving synthetic data while remaining computationally inexpensive.
@@ A.1 Variance of Learning Gradients in Diffusion Models
We show that the variance of the score function at an $n$-singular point increases asymptotically with respect to $n$ dimensions. Assume $x$ is a noisy observation from a Gaussian centered at a weighted one-hot vector $\alpha_{t}e_{k}\in\mathbb{R}^{K}$. We can define the forward diffusion process as:
@@
Hence, the score function of the conditional distribution $p_{t}(x|e_{k})$ is given by:
@@
$q(e_{k}|x)$ $=\frac{p(x|e_{k})p(e_{k})}{\sum_{m=1}^{K}p(x|e_{m})p(e_{m})}$ (21)
$=\frac{p(x|e_{k})}{\sum_{m=1}^{K}p(x|e_{m})}$ (22)
$=\frac{p(x|e_{k})}{\sum_{k\in S}p(x|e_{k})+\sum_{m\notin S}p(x|e_{m})}$ (23)
Since the forward process is modeled as: $p_{t}(x|e_{k})=\mathcal{N}(x|\alpha_{t}e_{k},\sigma_{t}^{2}I)$, we can infer that:
$p(x|e_{k})\propto\exp\left(-\frac{1}{2\sigma_{t}^{2}}||x-\alpha_{t}e_{k}||^{2}\right)$ (24)